how do you do find the perimeter of a triangle with the coordinates of: X(0,2),Y(4,-1),Z(-2,-1)
Answer by Maths68(1474)

Distance between two points
d= √(x2-x1)^2+(y2-y1)^2
X(0,2),Y(4,-1)
Distance between X and Y=a= √(4-0)^2+(-1-2)^2=√(4)^2+(-3)^2=√16+9=√25=5
Y(4,-1),Z(-2,-1)
Distance between Y and Z=b=√(-2-4)^2+(-1+1)^2=√(-6)^2+(0)^2=√36=6
X(0,2),Z(-2,-1)
Distance between Z and X=c=√(-2-0)^2+(-1-2)^2=√(-2)^2+(-3)^2==√4+9==√13=3.6
Primeter of triangle=a+b+c=5+6+3.6=14.6